OpenStudy (anonymous):

Anyone know what Let A = {1, 2, 3,4 }, and let R be a binary relation on A × A given by ((a,b),(c,d))∈R means? The A x A confuses me Anyone know what Let A = {1, 2, 3,4 }, and let R be a binary relation on A × A given by ((a,b),(c,d))∈R means? The A x A confuses me @Mathematics

OpenStudy (anonymous):

A x A is a cartesian product

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Binary_relation

myininaya (myininaya):

example: \[B=\{1,2\}\] say we want to find \[B \times B\] so \[B \times B=\{1,2\} \times \{1,2\}=\{(1,1),(1,2),(2,1),(2,2)\}\]

OpenStudy (anonymous):

Oh so all the possible combinations of them?

OpenStudy (anonymous):

((a,b),(c,d))∈R? And this just shows you what form you need to write it in?

OpenStudy (anonymous):

((a,b),(c,d))∈R? And this just shows you what form you need to write it in?

OpenStudy (anonymous):

the entire question actually also says and if and only if a divides c and b divides d, I don't see how a can ever divide see if we assume a is 1

myininaya (myininaya):

another example: \[C=\{a,b,c\} \] \[B=\{1,2\}\] \[C \times B=\{(a,1),(a,2),(b,1),(b,2),(c,1),(c,2)\}\] \[B \times C=\{(1,a),(2,a),(1,b),(2,b),(1,c),(2,c) \}\] do you see the difference?

OpenStudy (anonymous):

Oh yeah I see the difference

myininaya (myininaya):

well ok so how is R defined R is anything in the form ((a,b),(c,d)) if and only if a divides c and b divides d where (a,b) and (c,d) are elements of A cross A ?

myininaya (myininaya):

so we need to look in the set A cross A such that we can find two ordered pairs in that set that satisfy a|c and b|d

OpenStudy (anonymous):

Yeah I'm suppose to show it's an order

myininaya (myininaya):

example i will name two pairs of ordered pairs that is an element of R you should try to find the rest \[((2,1 ),(4,2)) \in R\] see 2|4 and 1|2

OpenStudy (anonymous):

but does 2/4?

myininaya (myininaya):

2 divides 4 is 2|4 which means there is some integer k such that 2k=4

OpenStudy (anonymous):

Oh I see okay so after I get all those pairs,I use them to find it the binary is reflexive, transitive and symmetric?

myininaya (myininaya):

\[a|c=> ak=c , k \in \mathbb{Z}^+\]

OpenStudy (anonymous):

Okay I see, but do you know what transitive means? sorry I keep asking so many questions you explain them really well

OpenStudy (anonymous):

I get that it's like x,y y,z so z,x is true or something but still kind of confused b/c what if I only have

OpenStudy (anonymous):

like (a,b) and [c,d)

OpenStudy (anonymous):

am I going to assume (a,b) (b,c) c,d so a,d?

myininaya (myininaya):

i kind of don't understand what you're asking

myininaya (myininaya):

are you suppose to just find all the elemetns of A crosss A or you trying to show R is a relation

OpenStudy (anonymous):

I'm suppose to show that R is an orde rso show that it's symmetric, reflexive and transitive but I don't really understand transitive

OpenStudy (anonymous):

and also doesn't everything divide with 1,1?

OpenStudy (anonymous):

so isn't it all elements?

myininaya (myininaya):

1 divides everything

OpenStudy (anonymous):

So I would have 16 elements?

OpenStudy (anonymous):

so would I have to go like I got (1,1) (1,2) (1,3) (1,4) (2,1) (2,2) (2,3) etc. so say 2,1 and 1,2 I'd get 2,2

OpenStudy (anonymous):

b/c (x,y) (y,z) therefore (x,z) right so 2,2 is in my set but would I have to do this test for all the elements?

myininaya (myininaya):

A = {1, 2, 3,4 } \[A \times A=\{1,2,3,4\} \times \{1,2,3,4\}\] \[=\{(1,1),(1,2),(1,3),(1,4),(2,1),(2,2),(2,3),(2,4),\] \[(3,1),(3,2),(3,3),(3,4),(4,1),(4,2),(4,3),(4,4)\}\]

myininaya (myininaya):

\[((1,1),(i,j)) \in R\] where i=1,2,3,4 and j=1,2,3,4 this is true since 1|i and 1|j

myininaya (myininaya):

\[((1,2),(i,j)) \in R\] where i=1,2,3,4 and j=4

myininaya (myininaya):

oops j=2,4

myininaya (myininaya):

since 2 divides itself

myininaya (myininaya):

\[((1,3),(i,j)) \in R\] where i=1,2,3,4 and j=3

myininaya (myininaya):

\[((1,4),(i,j)) \in R\] where i=1,2,3,4 and j=4

myininaya (myininaya):

\[((2,1),(i,j)) \in R\] where i=2,4 and j=1,2,3,4

myininaya (myininaya):

\[((2,2),(i,j)) \in R\] where i=2,4 and j=2,4

myininaya (myininaya):

\[((2,3),(i,j)) \in R\] where i=2,4 and j=3

myininaya (myininaya):

\[((2,4),(i,j)) \in R\] where i =2,4 and j=4

myininaya (myininaya):

\[((3,1),(i,j)) \in R\] where i=3 and j=1,2,3,4

myininaya (myininaya):

\[((3,2),(i,j)) \in R\] where i=3 and j=2,4 -------------------------- \[((3,3),(i,j)) \in R\] where i=3 and j=3 (i wonder can we use on element from A cross A twice) -------------------------- \[((3,4),(i,j)) \in R\] where i=3 and j-4 (i wonder...) -------------------------- \[((4,1),(i,j)) \in R\] where i=4 and j=1,2,3,4 -------------------------- \[((4,2),(i,j)) \in R\] where i=4 and j=2,4 -------------------------- \[((4,3),(i,j)) \in R\] where i=4 and j=3 ( i wonder ...)

OpenStudy (anonymous):

Oh and (4,4) is everything except anything with 3 in it

myininaya (myininaya):

we either have ((4,4),(4,4)) is in R or we don't say it is in R thats what all those wondering statements were about

myininaya (myininaya):

i think we do say it is in R

myininaya (myininaya):

maybe zarkon will say something

OpenStudy (anonymous):

okay but then what do I do those i,j pairs?

OpenStudy (anonymous):

do I now use those to find R is an order?

myininaya (myininaya):

oh i thought you were looking for all the elements of R

OpenStudy (anonymous):

Oh sorry if my question is not making sense. I feel bad for making you write all that out

OpenStudy (anonymous):

I'm just trying to show that R is an order

OpenStudy (anonymous):

Oh it says show so I assumed it does but it's alright I'll try to figure it out. Thanks for trying and spending time trying to explain it!! Really appreciate it

myininaya (myininaya):

for all x, y and z in X it holds that if xRy and yRz then xRz is the definition for transtive so we have that we need to show \[(r,s)R(t,u) \& (t,u)R(x,y)=> (r,s)R(x,y)\] lets see if this makes a difference with our example (1,1)R(2,4) and (2,4)R(4,4)=>(1,1)R(4,4) ok that works

myininaya (myininaya):

my early definition was alittle off

OpenStudy (anonymous):

so (1,1)R(4,4) needs to be in the set of elements?

OpenStudy (anonymous):

Sorry what does the R stand for when you write it like 1,1)R(4,4)

myininaya (myininaya):

so anyways we need to prove this so... if (r,s)R(t,u) and (t,u)R(x,y) then we have r|t and s|u and t|x and u|y we need to show (r,s)R(x,y)=> r|x and s|y --------------- so we have r|t and t|x this means ri=t and ta=x for some positive integers i and a so we are trying to show rp=x (ri)a=x r(ia)=x ia is an integer since i and a are integers and the set of integers in closed by multiplication so we have shown r|x now we nned to show s|y so we have s|u and u|y s|u=> sv=u and u|y=> ue=y for some positive integers v and e so we need to show s|y=> sq=y for some integer q (sv)e=u s(ve)=u ve is an integer since the set of integers is closed by multiplication so s|u we have now proved the transtive property for this relation R

myininaya (myininaya):

you asked what does xRy mean The statement (x,y) ∈ R is read "x is R-related to y", and is denoted by xRy or R(x,y).

OpenStudy (anonymous):

Oh so we don't need to show each element physically? I understand what you did I can do the same with reflexive and anti symmetric right

myininaya (myininaya):

yes

myininaya (myininaya):

you don't need to go through each element of R to see you can do one example for each property if you want thats what i did before i proved the transitive part

OpenStudy (anonymous):

Okay thanks so much for your help!! Wish I could give you more medals, really appreciate it thanks again!

myininaya (myininaya):

but it is not nessarcy

myininaya (myininaya):

ok